### v2

# A Bayesian Flow Network Framework for Chemistry Tasks

Abstract
In this work, we introduce ChemBFN, a language model that handles chemistry tasks based on Bayesian flow networks working on discrete data. A new accuracy schedule is proposed to improve the sampling quality by significantly reducing the reconstruction loss. We show evidence that our method is appropriate for generating molecules with satisfied diversity even when a smaller number of sampling steps is used. A classifier-free guidance method is adapted for conditional generation. It is also worthwhile to point out that after generative training, our model can be fine-tuned on regression and classification tasks with the state-of-the-art performance, which opens the gate of building all-in-one models in a single module style. Our model has been open sourced at https://github.com/Augus1999/bayesian-flow-network-for-chemistry.
keywords:
molecule generation,molecular property prediction,reaction yield prediction
[…]
1 Introduction
Autoregressive models (ARs) including SMILES-based or fragment-based models1, 2, 3, 4, 5, 6, 7, 8, 9 that leverage the power of language models (LMs) and reinforcement learning7, 8, 9 and graph-based models10, 11, 12, 13, 14, 15 coupled with advanced techniques such as Monte Carlo tree search11, 12, 13 have been proved their success in several de novo design benchmarks16, 6 consisted of drug-like molecules. The constraint of ARs, i.e., the number of sampling steps is the size of generated object, however, limits the potential of generating large molecules. Conversely, the recently emerging denoising-diffusion models17 (DMs) offer a way to generate objects of any size within a fixed sequence of sampling process. However, it has been pointed out in the research of C. Vignac et al18 that SMILES-based models generally worked better than graph DMs even when a dedicatedly designed discrete diffusion method was applied.
Bayesian flow networks19 (BFNs) are in a different category of generative models that decouple the sampling process with the size of generated objects as well. Different from DMs, BFNs directly work on the parameters of data distributions which naturally enable them to handle both continuous (including discretised) and discrete data without any data preprocessing or change of (mathematical) framework. Although the authors of BFN showed evidence in the original paper19 that BFN advantaged over discrete DMs on discrete data generating, e.g., text generation, the recent researches considering de novo molecule design only successfully employed it on continuous and discretised data, e.g., 3D molecular conformation generation20 rather than language-like representations such as SMILES21 or SELFIES22. One potential reason discouraging the application to text generation is the lack of exact analytical expression for the accuracy schedule $\beta(t)$, one critical component of BFNs, in the discrete case, while the speculated quadratic $\beta(t)$ in the original paper is, as admitted by the authors19, suboptimal.
In this paper, we introduce ChemBFN, a Bayesian Flow Network framework for Chemistry tasks, that leverages our newly proposed accuracy schedule and transformer23 encoder model to generate 1D language-like molecular representations e.g., SMILES and SELFIES. The experiments demonstrated that models with our accuracy schedule outperform those with the quadratic accuracy schedule. Besides, the generative training of BFN method can be a powerful pretraining strategy for downstream tasks in molecular property predictions, including regressions and classifications, and reaction yield predictions.
2 Methods
2.1 Discrete Bayesian Flow Networks
A functional BFN is consisted of a neural network (NN) model that converts the input distribution $\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})$ into the output distribution $\boldsymbol{p}_{O}(\boldsymbol{x}|\boldsymbol{\theta};t)$ and a Bayesian update process that updates the pervious input distribution to the current state according to a sender distribution $\boldsymbol{p}_{S}(\boldsymbol{y}|\boldsymbol{x};\alpha)$, where $\boldsymbol{\theta}$ is the parameter of data $\boldsymbol{x}$, and $\boldsymbol{y}$ is a sample of $\boldsymbol{x}$19. The none-negative monotonic increasing function $\alpha$, namely accuracy rate, guides the sender distribution to moving to a more informative direction along with the time19. Since $\alpha$ can be either continuous or discretised, a continuous accuracy schedule $\beta(t)$ is defined instead, which generates $\alpha$ as
$\alpha=\left\{\begin{array}[]{ll}{\dfrac{d}{dt}\beta(t)},&{\textrm{when $\alpha$ is continuous}}\\ {\beta(t_{i})-\beta(t_{i-1})},&{\textrm{when $\alpha$ is discretised}}.\end{array}\right.$ (1)
In the discrete case, we have (1) all the distributions are K-class categorical distributions; (2) the sample is defined as $\boldsymbol{y}=\mathcal{N}(\alpha(K\boldsymbol{e_{x}}-\boldsymbol{1}),\alpha K\boldsymbol{I})$ when a Gaussian sampling is utilised, where $\boldsymbol{e_{x}}$ is the one-hot representation of data $\boldsymbol{x}$; (3) the Bayesian update function is defined as $h(\theta^{(d)},y^{(d)},\alpha)=e^{y^{(d)}}\theta^{(d)}/\sum_{k=1}^{K}e^{y^{(d)}_{k}}\theta^{(d)}_{k}$, where $\cdot^{(d)}$ is the $d^{th}$ parameter19.
During the training stage, a receiver distribution $\boldsymbol{p}_{R}(\hat{\boldsymbol{y}}|\boldsymbol{\theta};t,\alpha)$ is drawn by sampling the output of NN with the same sampling method as sender distribution19. The model is optimised by minimising the Kullback-Leibler divergence between the receiver distribution and the sender distribution, which is decoupled as a $n$-step loss ($L^{n}$) and a reconstruction loss ($L^{r}$) and only the first loss in practice is used19. The limit case, i.e, continuous time loss, $L^{\infty}=\lim_{n\rightarrow\infty}L^{n}$ has been proved more efficient19. During the sampling (generating) stage, since the receiver distribution has been trained to match the sender distribution, i.e., $\boldsymbol{p}_{R}(\hat{\boldsymbol{y}}|\boldsymbol{\theta};t,\alpha)\sim\boldsymbol{p}_{S}(\boldsymbol{y}|\boldsymbol{x};\alpha)$, the receiver distribution is used in Bayesian update process directly to update the input distribution (initialised as a uniform distribution over K categories) where a discretised $\alpha$ is employed.
2.2 Model Architecture
Our model is an adaptation of DiT24 model. The differences in our implementation include (1) the use of categorical distributions rather than image embeddings for input tokens because we are not dealing with images; (2) logits output that are then transformed into probabilities by softmax function; (3) the replacement of activation function with SELU25 function; (4) the use of a 2-layer multilayer perceptron (MLP) to form the time embedding since “time” in BFN is continuous from 0 to 1; (5) the employment of XPOS26 variation of rotary positional embedding27. The architecture is shown in Figure 1.
Figure 1: Visualised scheme of our model. The architecture is inspired by DiT24. The multi-head self-attention layers did not use causal masking which is the same as BERT28 while we replaced the commonly used positional embedding method (absolute positional embedding used in DiT, BERT and RoBERTa29 models) with the novel XPOS26 variation of rotary positional embedding27. Note that each FFN (feed-forward network) layer adsorbs a dropout layer.
Following the notations of the BFN paper19, the parameter of categorical distributions inputted into the neural network is denoted by $\boldsymbol{\theta}=(\theta^{(1)},\theta^{(2)},...,\theta^{(D)})\in[0,1]^{KD}$ ($K$ is the number of categories, $D$ is the number of input data, and $\theta^{(d)}$ is the $d^{th}$ parameter) and the output distribution at time step $t$ is denoted by $\boldsymbol{p}_{O}(\cdot|\boldsymbol{\theta};t)\in[0,1]^{KD}$. We denote the sum of time embedding vector and conditioning vector as $\boldsymbol{c}$. A null conditioning $\phi$ is equivalent to a zero vector $\boldsymbol{0}$.
In each experiment described in the later text, we employed the same hyperparameters of the model except category number $K$ that depends on molecular representations. The 2-layer MLP with SELU activation has the shape of [1, 256, 512]. We employed 12 Transformer layers, of which had 8 attention heads each, with the attention temperature $\tau=\sqrt{2d_{h}}$ ($d_{h}$ is the feature number of each attention head)30. The dropout rate was 0.01 and the hidden feature number was 512. These settings lead to a total learnable parameters of the model of the magnitude of 54M.
2.3 A New Accuracy Schedule
In the case of BFN, an accuracy schedule function $\beta(t)$ drives the expectation of entropy of the input distribution $\mathbb{E}_{\boldsymbol{p}_{F}(\boldsymbol{\theta}|\boldsymbol{x};t)}H[\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})]$ to decrease linearly with $t$, where $\boldsymbol{x}$ stands for the clear data, $\boldsymbol{p}_{F}(\boldsymbol{\theta}|\boldsymbol{x};t)$ represents Bayesian flow distribution, and $\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})$ is the input distribution as denoted in the original paper19. The mathematical difficulty of deriving the expectation analytically in the discrete case compels us to speculate from intuition. The authors of BFN claimed that “$\beta(t)=t^{2}\beta(1)$ was a reasonable approximation”, but disclosed later that finding a suitable value for the hyperparameter $\beta(1)$ was not an easy job19.
Here, we give our estimation of $\beta(t)$. If we estimate the expected entropy of the input distribution (denoted as $E$ for short) as $E\sim f(K)e^{-\frac{K}{4}\beta(t)}$, then the relationship $E(t)=(1-t)E(0)+tE(1)$ that eliminates the unknown factor $f(K)$ gives us
$\beta(t)=-\frac{4}{K}\ln{\left(1-t+te^{-\frac{K}{4}\beta(1)}\right)}$ (2)
and the corresponding
$\alpha(t)=\frac{d\beta}{dt}=\frac{4}{K}\frac{1-e^{-\frac{K}{4}\beta(1)}}{1-t+te^{-\frac{K}{4}\beta(1)}},$ (3)
where $\beta(1)$ is still a hyperparameter. Equation (3) changes the continuous time loss $L^{\infty}$ to
$L^{\infty}(\boldsymbol{x})=\frac{K}{2}\mathbb{E}_{t\sim U(0,1),\boldsymbol{p}_{F}(\boldsymbol{\theta}|\boldsymbol{x};t)}\left(\alpha(t)\|\boldsymbol{e_{x}}-\hat{\boldsymbol{e}(\boldsymbol{\theta};t)}\|^{2}\right),$ (4)
where $\boldsymbol{e_{x}}$ is the one-hot representation of data $\boldsymbol{x}$ while $\hat{\boldsymbol{e}(\boldsymbol{\theta};t)}$ is the predicted categorical distribution of data $\boldsymbol{x}$ at time $t$. Note that when $\beta(1)$ is large, $\alpha(1)$ goes to extremely large. Therefore, we limit $\alpha(1)\leq 32\beta(1)$, from which
$\beta(1)_{max}\approx 20.4054/K$ (5)
is obtained. An example of how our accuracy schedule looks different from original one is plotted in Figure 2. We shall show in the later experiments that our $\beta(t)$ in Equation (2) works better than quadratic ones.
Figure 2: Comparing our accuracy schedule with quadratic accuracy schedule initialised with the same value of $\beta(1)$. (Left) Accuracy schedules $\beta(t)$. (Right) The accuracy rates $\alpha(t)$. Note that our $\beta(t)$ does not deviate too much from quadratic one, yet the rate (derivative) differs substantially as $t$ goes to 1.
2.4 Datasets and Benchmarks
Two benchmarks – MOSES16 and GuacaMol6 – were used to evaluate the generative performance, e.g., the similarity between generated molecules and training molecules, of ChemBFN. We reported the distribution-learning metrics of these benchmarks in Experiments and Results. A summary of these metrics is in Table 1.
Table 1: A brief summary of used metrics of MOSES and GuacaMol benchmarks
Metrics Description
Valid Fraction of valid molecules
Unique Fraction of unique molecules
IntDiv1 and IntDiv2 Internal diversities
Novelty Fraction of generated unseen molecules compared with training data
FCD Fréchet ChemNet Distance31
SNN Tanimoto similarity to a nearest neighbour
Frag BRICS fragment32 cosine similarity
Scaf Bemis–Murcko scaffold33 cosine similarity
Filter Fraction of molecules that fit pre-defined constructions
KL Divergence Kullback-Leibler divergence
The QM934 dataset was employed to study the capability of conditional generation of our method. We randomly selected 110k molecules, before which 3054 invalid data were removed, with the triple $(\epsilon_{HOMO},\epsilon_{LUMO},\Delta\epsilon_{HOMO-LUMO})$ as the conditioning label to form the training set.
In order to evaluate the downstream performance, 40M unique SMILES and 190M unique SMILES strings were randomly selected from easily accessed ZINC1535 database that formed two pretraining sets. The model trained on the 40M set was finetuned on several regression (ESOL, FreeSolv, Lipo, etc.) and classification (BBBP, BACE, ClinTox, HIV, etc.) tasks, including the subsets of widely used MoleculeNet36 benchmark. A brief description of used MoleculeNet tasks is in Table 2. Each dataset was split into training/validation/testing sets in the ratio of 80/10/10 following the scaffold splitting method proposed in DeepChem37 project. We reported ROC-AUC (area under receiver operating characteristic curve) for classification tasks and RMSE (root-mean squared error) for regression tasks in Experiments and Results. In addition to the tasks of MoleculeNet, two less biased datasets – the public ADME dataset published by C. Fang et al38 consisted of 6 dedicatedly collected absorption, distribution, metabolism, and excretion (ADME) in vitro endpoints together with a Kinase inhibitor dataset prepared by J.Wu et al39 that contains bioactivities of total 141,086 compounds for 354 kinases – were employed to further benchmark our method in activity prediction. A brief summary of sub-tasks of ADME dataset is in Table 2. For ADME dataset We employed the same split provided by C. Fang et al38; for Kinase inhibitor dataset, we prepared a random split and a scaffold split (both had training/validation/testing = 80/10/10). The testing MAE, RMSE, Pearson’s correlation coefficient (R value), and averaged ROC-AUC were reported in Experiments and Results.
Table 2: A brief summary of used MoleculeNet and public ADME tasks
Name № molecules № tasks Label
ESOL 1,128 1 Aqueous solubility $log_{10}(S/{\rm mol\cdot L^{-1}})$
FreeSolv 642 1 Experimental hydration free energy / kcal/mol
Lipo 4,200 1 Octanol/water distribution coefficient $log_{10}D_{7.4}$
HLM 3,087 1 Logarithm of human liver microsomal stability / $\rm mL\cdot min^{-1}\cdot kg^{-1}$
RLM 3,054 1 Logarithm of rat liver microsomal stability / $\rm mL\cdot min^{-1}\cdot kg^{-1}$
hPPB 1,808 1 Logarithm of human plasma protein binding (percent unbound)
rPPB 884 1 Logarithm of rat plasma protein binding (percent unbound)
[…]
Solubility 2,173 1 Aqueous solubility $log_{10}(S/{\rm\mu g\cdot mL^{-1}})$ at PH = 6.8
BBBP 2,039 1 If a compound penetrates the blood-brain barrier
BACE 1,513 1 If a compound inhibits BACE-1 protein
ClinTox 1,478 2 Task 1: if a drug has been approved by FDA
Task 2: if the same drug had toxicity during clinical trail
HIV 41,127 1 If a compound is an HIV inhibitor
The USPTO-50k40 dataset, Buchwald-Hartwig and Suzuki-Miyaura reaction yield datasets from high-throughput experiments (HTE) cleaned by P. Schwaller et al41 were employed to train the model to predict reaction yield. USPTO-50k that contains 50k reactions mined from patents were used to pre-train the model while HTE data were used for fine-tuning. We report coefficient of determination (R2 score) on testing sets in Experiments and Results.
AqSolDB42, a more challenging solubility dataset containing more species than ESOL, was used to investigate the effect of the size of pretraining data. A training/validation/testing (80/10/10) split was generated using scaffold splitting method. Testing MAE (mean absolute error) and RMSE were reported in Experiments and Results.
[…]
2.5 Fine-tuning Strategy
[…]
3 Experiments and Results
3.1 Unconditional Generation
Figure 4: Visualisation of the impact on training loss, reconstruction loss $L^{r}$ and continuous (cts) time loss $L^{\infty}$ of different accuracy schedules with different values of $\beta(1)$. $L^{r}$ and $L^{\infty}$ were computed on 1k discretised steps after training.
We first evaluate the effect of different $\beta(t)$ with different values of $\beta(1)$ using MOSES dataset. We reported the validity, FCD on scaffold set, SNN on scaffold set, Frag on scaffold set, Scaf on scaffold set, Filters, and Novelty scores computed by MOSES program in Table 3 together with reconstruction loss $L^{r}=-\mathbb{E}_{\boldsymbol{p}_{F}(\boldsymbol{\theta}|\mathbf{x};t)}\ln{\boldsymbol{p}_{O}(\mathbf{x}|\boldsymbol{\theta};t)}$ and continuous time loss $L^{\infty}$ in Figure 4. It is clear that raising $\beta(1)$ in both quadratic and our schedules did not have obvious influence on training loss but lowered $L^{r}$, while our schedule lead to a lower loss when $\beta(1)$ was the same. The effect on $L^{\infty}$ was subtle. However, after we calculated the $R^{2}$ values of the cumulative $L^{\infty}$ curves, we found that while using quadratic $\beta(t)$ the curve became more distorted when $\beta(1)$ was larger ($R^{2}|_{\beta(1)=0.0448}=0.995$ while $R^{2}|_{\beta(1)=0.15}=0.992$); After switching to our $\beta(t)$ the curves were more linear (i.e., $L^{\infty}$ was more uniform) and the linearity was not affected by the value of $\beta(1)$ ($R^{2}|_{\beta(1)=0.0448}=R^{2}|_{\beta(1)=0.0829}=0.997$). The metrics in Table 3 provide more quantitative evidences that our $\beta(t)$ is more optimal. It is notable that a larger $\beta(1)$ value usually result to better scores. Therefore, we conclude here that our proposed $\beta(t)$ with $\beta(1)=\beta(1)_{max}=20.4054/K$ is a more optimal solution in discrete BFNs.
Table 3: Comparing scores of MOSES benchmark when varying $\beta(1)$ value of different accuracy schedulesa
[…]
[][c] $\beta(1)$ Valid $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Filters $\uparrow$ Novelty $\uparrow$ quad 0.15 0.893 $\pm$ 0.001 3.438 $\pm$ 0.034 0.559 $\pm$ 0.000 0.985 $\pm$ 0.000 0.095 $\pm$ 0.001 0.982 $\pm$ 0.000 0.900 $\pm$ 0.002 0.0829 0.895 $\pm$ 0.001 3.772 $\pm$ 0.012 0.551 $\pm$ 0.001 0.984 $\pm$ 0.001 0.096 $\pm$ 0.006 0.985 $\pm$ 0.001 0.900 $\pm$ 0.002 0.0448 0.899 $\pm$ 0.003 3.902 $\pm$ 0.045 0.561 $\pm$ 0.000 0.988 $\pm$ 0.000 0.089 $\pm$ 0.006 0.986 $\pm$ 0.001 0.887 $\pm$ 0.003 ours 0.0829 0.900 $\pm$ 0.001 2.731 $\pm$ 0.015 0.563 $\pm$ 0.000 0.990 $\pm$ 0.000 0.091 $\pm$ 0.004 0.987 $\pm$ 0.001 0.886 $\pm$ 0.000 0.0448 0.900 $\pm$ 0.001 3.580 $\pm$ 0.008 0.568 $\pm$ 0.000 0.987 $\pm$ 0.000 0.075 $\pm$ 0.006 0.987 $\pm$ 0.000 0.877 $\pm$ 0.001
  • a
    $\uparrow$ indicates that the higher is better and $\downarrow$ stands for the contrary. The best results are in bold. We used a sampling step of 1k.
In the above experiments, we used a dynamic padding strategy, i.e., each batch were padded to the maximum length of that batch, to reduce the training time. In the following experiments, global padding strategy, i.e., padding all batches to a global maximum length, was employed to compare with dynamic strategy on both MOSES and GuacaMol benchmarks. The results were summarised in Table 4. We found that the global padding method benefited the performance. In the following experiment, we therefore employed the global padding method in generative tasks.
Table 4: Scores of MOSES and GuacaMol benchmarks when different padding strategies were used during traininga
[…]
[][c] Strategy MOSES Valid $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Filters $\uparrow$ Novelty $\uparrow$ dynamic 0.900 $\pm$ 0.001 2.731 $\pm$ 0.015 0.563 $\pm$ 0.000 0.990 $\pm$ 0.000 0.091 $\pm$ 0.004 0.987 $\pm$ 0.001 0.886 $\pm$ 0.000 global 0.916 $\pm$ 0.001 2.730 $\pm$ 0.014 0.565 $\pm$ 0.001 0.990 $\pm$ 0.000 0.094 $\pm$ 0.002 0.987 $\pm$ 0.001 0.880 $\pm$ 0.002 GuacaMol Valid $\uparrow$ Unique $\uparrow$ Novelty $\uparrow$ KL Divergence $\uparrow$ FCD $\uparrow$ dynamic 0.799 $\pm$ 0.003 0.815 $\pm$ 0.002 0.975 $\pm$ 0.000 0.810 $\pm$ 0.001 0.370 $\pm$ 0.003 global 0.807 $\pm$ 0.003 0.818 $\pm$ 0.001 0.975 $\pm$ 0.001 0.808 $\pm$ 0.010 0.399 $\pm$ 0.002
  • a
    $\uparrow$ for higher is better and $\downarrow$ for contrary. The best results are in bold. We used a sampling step of 1k.
Finally, we trained models applying the above optimal settings (i.e., $\beta(1)=20.4054/K$ and global padding) on MOSES and GuacaMol datasets. Both SMILES and SELFIES versions were implemented. The comparison with published state-of-the-art (SOTA) models5, 3, 10, 4, 12, 18, 6 are summarised in Table 5, Table 6, and Table 7. We found that (1) except FCD, metrics of both SMILES version and SELFIES version were close to SOTA performance. (2) number of sampling step as expected affected the validity of generated molecules (for SMILES version only because SELFIES always gives valid molecules22), but dropping from 1k steps to 100 steps did not degrade the performance a lot. If lower validity is acceptable, only sampling 10 steps significantly reduce the computational time without much impact on other qualities. Larger FCD (in the term of GuacaMol is lower FCD score where FCD score = $e^{-0.2\rm FCD}$) is a hint that BFNs learn the grammar of molecules rather than the way of combining characters within the dataset.
Table 5: Testing metrics on MOSES test set compared with SOTA modelsa
Method Valid $\uparrow$ Unique@1k $\uparrow$ Unique@10k $\uparrow$ IntDiv1 $\uparrow$ IntDiv2 $\uparrow$ Novelty $\uparrow$
ARs JTN-VAE5 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 0.855 $\pm$ 0.003 0.850 $\pm$ 0.004 0.913 $\pm$ 0.006
LatentGAN3 0.897 $\pm$ 0.003 1.0 $\pm$ 0.0 0.997 $\pm$ 0.000 0.857 $\pm$ 0.001 0.851 $\pm$ 0.001 0.950 $\pm$ 0.001
GraphINVENT10 0.964 1.0 0.998 0.857 0.851
MolGPT4 0.994 1.0 0.857 0.851 0.797
DMs DiGress18 0.857 1.0 0.950
BFNs ChemBFN10 0.835 $\pm$ 0.003 1.0 $\pm$ 0.0 0.999 $\pm$ 0.000 0.851 $\pm$ 0.000 0.844 $\pm$ 0.000 0.921 $\pm$ 0.002
ChemBFN100 0.911 $\pm$ 0.002 1.0 $\pm$ 0.0 0.998 $\pm$ 0.000 0.837 $\pm$ 0.000 0.831 $\pm$ 0.000 0.884 $\pm$ 0.002
ChemBFN1k 0.916 $\pm$ 0.001 1.0 $\pm$ 0.0 0.998 $\pm$ 0.000 0.836 $\pm$ 0.000 0.830 $\pm$ 0.000 0.880 $\pm$ 0.002
[…]
  • a
    The metrics of all other models were copied from the original paper. $\uparrow$ for the higher is better. (10, 100, 1k) are the number of sampling steps. * for SELFIES version. The best results are in bold.
Table 6: Metrics on MOSES scaffold test seta
Method FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaff $\uparrow$ Filters $\uparrow$
ARs JTN-VAE5 0.938 $\pm$ 0.053 0.519 $\pm$ 0.007 0.995 $\pm$ 0.000 0.101 $\pm$ 0.011 0.976 $\pm$ 0.002
LatentGAN3 0.828 $\pm$ 0.012 0.513 $\pm$ 0.000 0.997 $\pm$ 0.001 0.107 $\pm$ 0.010 0.974 $\pm$ 0.001
GraphINVENT10 1.223 0.539 0.986 0.127 0.950
DMs DiGress18 1.19 0.52 0.148 0.971
BFNs ChemBFN10 2.768 $\pm$ 0.035 0.533 $\pm$ 0.000 0.988 $\pm$ 0.000 0.145 $\pm$ 0.004 0.976 $\pm$ 0.001
ChemBFN100 2.604 $\pm$ 0.040 0.562 $\pm$ 0.001 0.991 $\pm$ 0.000 0.103 $\pm$ 0.005 0.985 $\pm$ 0.001
ChemBFN1k 2.730 $\pm$ 0.014 0.565 $\pm$ 0.001 0.990 $\pm$ 0.000 0.094 $\pm$ 0.002 0.987 $\pm$ 0.001
[…]
  • a
    Settings are the same as Table 5 while $\downarrow$ for the lower is better.
Table 7: Testing metrics on GuacaMol distribution-learning tasksa
Method Valid $\uparrow$ Unique $\uparrow$ Novelty $\uparrow$ KL Divergence $\uparrow$ FCD $\uparrow$
ARs MolGPT4 0.981 0.998 1.0 0.992 0.907
SMILES LSTM6 0.959 1.0 0.912 0.991 0.913
VGAE-MCTS12 1.0 1.0 1.0 0.659 0.009
DMs DiGress18 0.852 1.0 0.999 0.929 0.680
BFNs ChemBFN1k 0.807 $\pm$ 0.003 0.818 $\pm$ 0.001 0.975 $\pm$ 0.001 0.808 $\pm$ 0.010 0.399 $\pm$ 0.002
[…]
  • a
    Settings are the same as Table 5.
3.2 Conditional Generation of Small Molecules
The classifier-free guidance45 method is easily adapted into BFN, where only the computation of output distribution needs changing during sampling process. The pseudocode for computing discrete output distribution is presented in Algorithm 1. In the experiment, we jointly trained a model conditionally and unconditionally on QM9 dataset with an unconditional rate $p_{uncond}=0.2$. In the sampling stage, $w$ was set to 4. We sampled 10 molecules using the label [-0.249, 0.0615, 0.3105] that was transformed to $\boldsymbol{y}$ via a trained 2-layer MLP. 10 unconditioned samples were generated as a control group. RDKit46 was employed to generate the 3D conformations, then the geometry optimisations and energy calculations were performed via PySCF47 at B3LYP/6-31G(2df,p) level of accuracy. The results of MAE between calculated values and labels are presented in Table 8. The conditioned samples are displayed in Figure 5.
Algorithm 1 Invoking classifier-free guidance into output distribution
$w\in\mathbb{R}$, conditioning vector $\boldsymbol{y}$
function DISCRETE_OUTPUT_DISTRIBUTION($\boldsymbol{\theta}\in[0,1]^{KD}$, $t\in[0,1]$, $\boldsymbol{y}\in\mathbb{R}^{f}$)
     Input ($\boldsymbol{\theta}$, $t$, $\boldsymbol{y}$) to network, receive $\boldsymbol{\Psi}(\boldsymbol{\theta},t,\boldsymbol{y})$ as output
     if in training stage or $\boldsymbol{y}$ is $\phi$ then
         $\boldsymbol{p}_{O}(\cdot|\boldsymbol{\theta};t)\leftarrow{\rm softmax}(\boldsymbol{\Psi}(\boldsymbol{\theta},t,\boldsymbol{y}))_{dim=-1}$
     else
         Input ($\boldsymbol{\theta}$, $t$, $\phi$) to network, receive $\boldsymbol{\Psi}(\boldsymbol{\theta},t,\phi)$ as output
         $\boldsymbol{p}_{O}(\cdot|\boldsymbol{\theta};t)\leftarrow{\rm softmax}((1+w)\boldsymbol{\Psi}(\boldsymbol{\theta},t,\boldsymbol{y})-w\boldsymbol{\Psi}(\boldsymbol{\theta},t,\phi))_{dim=-1}$
     end if
     return $\boldsymbol{p}_{O}(\cdot|\boldsymbol{\theta};t)$
end function
Table 8: MAE on QM9 dataset w/ and w/o classifier-free guidance generationa
$\epsilon_{HOMO}$ / a.u. $\epsilon_{LUMO}$ / a.u. $\Delta\epsilon$ / a.u.
Conditional 0.00724 0.00981 0.01329
Unconditional 0.01901 0.04076 0.04104
  • a
    Smaller errors are in bold.
Figure 5: Conditioned samples on QM9. The number of sampling steps was 1k. Since QM9 exhaustively included stable small molecules made up of CHONF, only 4 conditioned samples and 5 unconditioned samples are novel.
3.3 Molecular Scaffold Extension
Here, we show a simple inpaint strategy can extend molecular scaffolds by using ChemBFN. In every sampling steps, parameters of input distributions are modified as $\boldsymbol{\theta}\leftarrow\boldsymbol{M}\odot\boldsymbol{e_{x}}+(1-\boldsymbol{M})\odot\boldsymbol{\theta}$ before being inputted into the network, where $\boldsymbol{M}$ is the mask and $\boldsymbol{e_{x}}$ is the one-hot representation of scaffold. Figure 6 shows an example of extending scaffold ‘Cc1cc(OC5)cc(C6)c1.’ by a model trained on MOSES SAFE48 version, a variation of SMILES. We found that inpainting sampling for 10 to 100 steps was sufficient to generate complex molecules.
Figure 6: An example of extended molecular scaffold. The scaffold is highlighted in red.
3.4 Finetuning on Prediction Tasks
In this section, we compare our model with SOTA models49, 50, 51, 52, 53, 43, 44, 54, 55, including graph-based and language-based which could be further classified as smaller scale natural language processing models (NLPs) and large language models (LLMs), on subsets of MoleculeNet benchmark. As shown in Table 9, our method outperformed SOTA language models on several tasks, especially ClinTox and BBBP. It is notable that ChemBERTa43 and ChemBERTa-244, which had a similar model size with ours, were pretrained on 77M molecules but had worse scores on 3 out of 5 tasks than ours. This indicated that BFN-style generative pretraining is a better strategy than masked language modeling and multitask regression pretraining. A similar observation applied to CaRRoBERTa model that coupled the knowledge of ChatGPT56 (which is far larger in scale than ours and is believed to have seen more chemical texts) and the distillation capability of RoBERTa29 method: our model outperformed CaRRoBERTa on 4 out of 5 tasks. However, when comparing with graph neural networks (GNNs) our model performed averagely 1.7% worse, especially on regression tasks.
Table 9: Testing metrics on sub-tasks of MoleculeNet benchmark with scaffold splitting compared with SOTA modelsa
[…]
[][c] Method ROC-AUC $\uparrow$ RMSE $\downarrow$ ClinTox BBBP BACE HIV ESOL FreeSolv Lipo GNNs Uni-Mol49 91.9 $\pm$ 1.8 72.9 $\pm$ 0.6 85.7 $\pm$ 0.2 80.8 $\pm$ 0.3 0.788 $\pm$ 0.029 1.480 $\pm$ 0.048 0.603 $\pm$ 0.010 MolKD50 83.8 $\pm$ 3.1 74.8 $\pm$ 2.3 80.1 $\pm$ 0.8 74.9 $\pm$ 1.7 GEM51 90.1 $\pm$ 1.3 72.4 $\pm$ 0.4 85.6 $\pm$ 1.1 80.6 $\pm$ 0.9 0.798 $\pm$ 0.029 1.877 $\pm$ 0.094 0.660 $\pm$ 0.008 Mole-BERT52 78.9 $\pm$ 3.0 71.9 $\pm$ 1.6 80.8 $\pm$ 1.4 78.2 $\pm$ 0.8 1.015 $\pm$ 0.030 0.676 $\pm$ 0.017 LLMs CaRRoBERTa53 84.16 $\pm$ 17.63 81.99 $\pm$ 4.19 80.73 $\pm$ 1.42 0.96 $\pm$ 0.09 1.02 $\pm$ 0.06 NLPs ChemBERTa43 73.3 64.3 62.2 ChemBERTa-244 60.1 74.2 79.9 0.744 AGBT55 76.3 SMILES Transformer54 70.4 70.1 72.9 ChemBFN (ours) 99.18 $\pm$ 1.77 95.74 $\pm$ 0.70 73.56 $\pm$ 1.22 79.37 $\pm$ 1.66 0.884 $\pm$ 0.003 1.418 $\pm$ 0.067 0.746 $\pm$ 0.001 $\Delta_{GNNs_{best}}$ +8% +28% -14% -2% +12% -4% +24% $\Delta_{LMs_{best}}$ +18% +17% -9% +9% -8% 0%
  • a
    The metrics of all other models were copied from their original paper. $\uparrow$ indicates that the higher is better and $\downarrow$ stands for the contrary. The best results are in bold. The best results within the same category (graph-based or language-based) are underlined. Percentages in the last two rows show the performance changes w.r.t the best models and the colour represents whether our model was better (in red) or not (in blue).
We further benchmarked our model on the public ADME dataset38 (regression task) and Kinase inhibitor dataset39 (classification task). The results for the public ADME dataset were summarised in Table 10. For the Kinase inhibitor dataset, the averaged ROC-AUC over 354 assays tested on the random split was (87.93 $\pm$ 14.05)and the averaged ROC-AUC tested on the scaffold split was (79.35 $\pm$ 18.71)%.
Table 10: MAE, RMSE, and Pearson’s correlation coefficient on the public ADME dataset.
HLM RLM hPPB rPPB MDR1-MDCK ER Solubility
MAE 0.359 $\pm$ 0.005 0.428 $\pm$ 0.001 0.365 $\pm$ 0.006 0.408 $\pm$ 0.006 0.337 $\pm$ 0.003 0.411 $\pm$ 0.008
RMSE 0.474 $\pm$ 0.004 0.556 $\pm$ 0.005 0.479 $\pm$ 0.012 0.549 $\pm$ 0.009 0.466 $\pm$ 0.007 0.630 $\pm$ 0.013
R 0.653 $\pm$ 0.002 0.685 $\pm$ 0.007 0.771 $\pm$ 0.001 0.700 $\pm$ 0.003 0.765 $\pm$ 0.007 0.588 $\pm$ 0.003
3.5 Reaction Yield Prediction
In order to predict the reaction yield, we first trained the generative model to understand chemical reaction by learning to predict the products. We developed an in-context style guidance that during training stage, only the parameters of product in reaction SMILES were predicted. This was achieved by always masking the input distribution of reactant/reagent and >>much-greater-than$>$ > tokens that were converted to the corresponding one-hot representation, i.e., $\boldsymbol{\theta}\leftarrow\boldsymbol{M}_{rr}\odot\boldsymbol{e_{x}}+(1-\boldsymbol{M}_{rr})\odot\boldsymbol{\theta}$, where $\boldsymbol{M}_{rr}$ is the mask for reactant, reagent, and >>much-greater-than$>$ > token.
The generative model was first pre-trained on USPTO-50k dataset then post-trained on Buchwald-Hartwig and Suzuki-Miyaura coupling datasets before the whole prediction model was fine-tuned. The testing scores compared with previous researches57, 41, 58 were reported in Table 11. It is notable that the Yield-BERT series41, 58 were based on a pre-trained RXNFP59 model which had been pre-trained on over 2M reactions while our model was pre-trained on 50k reactions. Despite the disadvantage of limited access of pretraining data, the performance of our method was still close to that of largely pretrained model on random-split sets and significantly better on out-of-sample predictions.
Table 11: R2 scores on different testing sets of HTE Buchwald-Hartwig and Suzuki-Miyaura reaction datasetsa
Method
Dataset Split MFF57 Yield-BERT41 Yield-BERT-DA58 ChemBFN (ours)
Buchwald- Hartwig Rand 70/30 0.927 $\pm$ 0.007 0.951 $\pm$ 0.005 0.969 $\pm$ 0.004 0.952 $\pm$ 0.008
Test 1 0.85 0.84 $\pm$ 0.010 0.82 $\pm$ 0.01 0.844 $\pm$ 0.002
Test 2 0.71 0.84 $\pm$ 0.03 0.90 $\pm$ 0.01 0.910 $\pm$ 0.001
Test 3 0.64 0.75 $\pm$ 0.04 0.63 $\pm$ 0.05 0.787 $\pm$ 0.034
Test 4 0.18 0.49 $\pm$ 0.05 0.43 $\pm$ 0.07 0.633 $\pm$ 0.082
Avg. 1-4 0.60 0.73 $\pm$ 0.15 0.69 $\pm$ 0.19 0.794 $\pm$ 0.118
Suzuki- Miyaura Rand 70/30 0.81 $\pm$ 0.02 0.796 $\pm$ 0.011
  • a
    The scores of all other models were copied from the original paper. The best results are in bold. The score of “rand 70/30” split was the 10-fold average value. Test 1-4 were out-of-sample splits.
3.6 Is Larger Pretrain Dataset Better?
We have seen that our model, although was pretrained on 40M molecules, outperformed the models pretrained on larger dataset on several prediction tasks. Here rises a question: does a larger pretraining dataset benefit our method? To answer this, three models were trained on AqSolDB dataset, of which one was trained from scratch, one was pretrained on 40M molecules from ZINC15 database, and the third one was pretrained on 190M molecules from ZINC15. We summarised the testing results in Table 12. Interestingly, the errors did not shrink when the pretraining data grew from 40M to 190M. However, compared with zero pretraining, an improvement in performance of $\geq$12.5can be confirmed.
Table 12: Testing metrics of models with different pretrain data sizes (0, 40M, and 190M) on AqSolDB dataset
From scratch Pretrained on 40M Pretrained on 190M
MAE 0.978 $\pm$ 0.016 0.837 $\pm$ 0.005 0.851 $\pm$ 0.021
RMSE 1.309 $\pm$ 0.014 1.131 $\pm$ 0.008 1.145 $\pm$ 0.034
3.7 Training Details
For all generative tasks, the models were trained for 100 epochs with the batch-size of 120 molecule/batch. The learning rate ($lr$) was $5.0\times 10^{-5}$ that was linearly increased (warm-up) from $10^{-8}$ during the first 1,000 training steps.
We pre-trained one model on 40M SMILES for 15 epochs with the batch-size of 512 on single A100 GPU and one model on 190M SMILES for 5 epochs with the effective batch-size of 1,024 ($2\times 512$) on 2$\times$×A100 GPUs. The warm-up strategy and $lr$ were the same as mentioned above.
During fine-tuning stages, models were trained for 100 epochs on labelled datasets. The batch-size, both for training and validation, was 32 on MoleculeNet benchmark, AqSolDB dataset, public ADME dataset, and Kinase inhibitor dataset; the training batch-size was 16 for reaction yield prediction. $lr_{max}$ was $10^{-4}$ that was warmed up from $10^{-7}$ during the first 1,000 steps for regression tasks and 100 steps for classification tasks. After warm-up stage, $lr$ decreased by 0.2 after the validation metrics stopped improving for 20 epochs unless the learning rate had reached $10^{-6}$. The dropout rate of prediction MLP head was fine-tuned for each case and we recommend to try from $\{0.0,0.1,0.5,0.7\}$. The validation metrics for regression and classification tasks were MAE and inverted accuracy (i.e., 1 - accuracy), respectively.
We employed AdamW60 with default hyperparameters implemented in PyTorch61 as the optimizer for all tasks.
4 Conclusion
ChemBFN, a Bayesian flow network framework for chemistry tasks of both generation and prediction, was developed in this work. The new accuracy schedule helped ChemBFN achieve competitive performance of discrete diffusion models and autoregressive models on generating large molecules. We proposed a BFN-style generative pretraining strategy that surpassed existing language-based transformer models on several classification and regression tasks. We believe this work provides a tool that can accelerate researches of both drug designing and filtering and give in helpful information for synthesis planning. However, we still leave gaps between graph-based models in prediction tasks, which we shall keep for the future research.
5 Data and Software Availability
The code, pre-trained models, and instructions necessary to reproduce the results of this study are available for download at https://github.com/Augus1999/bayesian-flow-network-for-chemistry.
6 Acknowledgements
We express our gratitude to the Research Center for Computational Science (RCCS) in Okazaki, Japan, and its maintenance team for providing computing resources, including A100 GPUs. This work is under project 24-IMS-C043 of RCCS. We also thank Dr. Maho Nakata who kindly lent us his own RTX 3080 GPU and Prof. Kazumasa Okada for discussion.
7 Conflict of Interest
There is no conflict of interest.
8 Funding Sources
The authors claim that there is no funding related to this research.